LIST OF PARTICIPANTS

Balashova Galina Sergeevna Department of Mathematics, Moscow Power Engineering Institute, Krasnokazarmennaya 14, 111250 Moscow, Russia

E-mail: lesha@ipng.msk.su

Bourdaud Gérard Université Paris 6, U.R.A. 754, Tour 46-0, Boîte 186, 4, place Jussieu, F-75252 Paris Cédex 05, France

E-mail: bourdaud@mathp7.jussieu.fr

Burenkov Viktor Ivanovich Department of Mathematics, College of Cardiff, University of Wales, 23 Senghennydd Rd, Cardiff CF2 4YH, United Kingdom

E-mail: burenkov@cardiff.ac.uk

de la Torre Alberto Departamento de Análisis Matemático, Facultad de Ciencias, Universidad de Málaga, 29071 Málaga, Spain

E-mail: torre r@ccuma.sci.uma.es

Evans W. D. Department of Mathematics, College of Cardiff, University of Wales, 23 Senghennydd Rd, Cardiff CF2 4YH, United Kingdom

E-mail: w.d.evans@cardiff.ac.uk

Gao Ji Department of Mathematics, Community College of Philadelphia, 1700 Spring Garden Street, Philadelphia, PA 19130-3991, U.S.A.

E-mail: ccp math@shrsys.hslc.org

Gogatishvili Amiran A. Razmadze Mathematical Institute, Z. Rukhadze st. 1, 380093 Tbilisi, Republic of Georgia

E-mail: gogati@imath.kheta.ge

Gol’dman Mikhail Krupskoi st. 8-1-187, 117311 Moscow, Russia

E-mail: seulydia@glas.apc.org

Gurka Petr Department of Mathematics, Technical Faculty, Czech Agricultural University, Kamýcká 129, 16521 Praha, Czech Republic

E-mail: gurka@tech.vsz.cz

Haroske Dorothee Mathematisches Institut, Fakultät für Mathematik und Informatik, Friedrich-Schiller-Universität, Leutragraben 1, 07743 Jena, Germany

E-mail: haroske@minet.uni-jena.de

Harris Desmod J. Pryske Cottage, Llanishen, Chepstow, Gwent NP6 6QD, United Kingdom

E-mail: w.d.evans@cardiff.ac.uk

Hoh Walter SFB 343, Universität Bielefeld, Postfach 100131, 33501 Bielefeld, Germany

E-mail: hoh@mathematik.uni-bielefeld.de

Jacob Niels Mathematisches Institut, Universität Erlangen-Nürnberg,

Bismarckstraße 1\frac{1}{2}, 91054 Erlangen, Germany Kalyabin Gennadii A. Samara State Aerospace University, Moskovskoye

shosse 34, 443096 Samara, Russia

E-mail: feliks@sgau.volgacom.samara.su

Karapetyants Nikolai Department of Mathematics, Rostov State University, ul. Zorge 5, 344104 Rostov-on-Don, Russia

E-mail: nkarapet@ns.unird.ac.ru

Kokilashvili Vakhtang A. Razmadze Mathematical Institute, Z. Rukhadze st. 1, 380093 Tbilisi, Republic of Georgia

E-mail: kokil@imath.kheta.ge

Kondratyuk Andrii Andriovich Artema st. 59/14,290053 Lviv, Ukraine

E-mail: kond@franko.lviv.ua

Koubová Blanka Mathematical Institute, Academy of Sciences of the Czech Republic, Žitná 25, 11567 Praha, Czech Republic

E-mail: pick@earn.cvut.cz

Kováčik Ondrej Department of Mathematics, Faculty of Civil Engineering, University of Transport and Communications, J. M. Hurbana 15, 01026 Žilina, Slovak Republic

E-mail: matemsv@uvt.utc.sk

Krbec Miroslav Mathematical Institute, Academy of Sciences of the Czech Republic, Žitná 25, 11567 Praha, Czech Republic

E-mail: krbecm@earn.cvut.cz

Kufner Alois Mathematical Institute, Academy of Sciences of the Czech Republic, Žitná 25, 11567 Praha, Czech Republic

E-mail: kufner@earn.cvut.cz

Lang Jan Mathematical Institute, Academy of Sciences of the Czech Republic, Žitná 25, 11567 Praha, Czech Republic

E-mail: lang@karlin.mff.cuni.cz

Leopold Hans-Gerd Mathematisches Institut, Fakultät für Mathematik und Informatik, Friedrich-Schiller-Universität, Leutragraben 1, 07743 Jena, Germany

E-mail: leopold@minet.uni-jena.de

Lin Lai-Jiu Department of Mathematics, National Changhua University of Education, Changhua, Taiwan

E-mail: maljlin@math.ncue.edu.tw

Malarski Mircea Mathematisches Institut, Fakultät für Mathematik und Informatik, Friedrich-Schiller-Universität, Leutragraben 1, 07743 Jena, Germany

E-mail: malarski@minet.uni-jena.de

Muscalu Camil Institute of Mathematics of the Romanian Academy, P.O. Box 1-764, RO-70700 Bucharest, Romania

E-mail: muscalu@roimar.imar.ro

Mustonen Vesa Department of Mathematical Sciences, University of Oulu, Linnanmaa, P.O. Box 333, 90570 Oulu, Finland

E-mail: vmustone@cc.oulu.fi

Nekvinda Aleš Department of Mathematics, Faculty of Civil Engineering,

Czech Technical University, Technická 2, Praha 6, Czech Republic Opic Bohumír Mathematical Institute, Academy of Sciences of the Czech

Republic, Žitná 25, 11567 Praha, Czech Republic

E-mail: opic@earn.cvut.cz

Pérez Carlos Departamento de Matemáticas, Universidad Autónoma de Madrid, 28049 Madrid, Spain

E-mail: cperezmo@ccuam3.sdi.uam.es

Pflüger Klaus Freie Universität Berlin, Institut für Mathematik I, Arnimallee 2-6, D 14195 Berlin, Germany

E-mail: pflueger@math.fu-berlin.de

Pick Luboš Mathematical Institute, Academy of Sciences of the Czech Republic, Žitná 25, 11567 Praha, Czech Republic

E-mail: pick@earn.cvut.cz

Rákosník Jiří Mathematical Institute, Academy of Sciences of the Czech Republic, Žitná 25, 11567 Praha, Czech Republic

E-mail: rakosnik@earn.cvut.cz

Rodriguez Christina Mathematisches Institut, Fakultät für Mathematik und Informatik, Friedrich-Schiller-Universität, Leutragraben 1, 07743 Jena, Germany

E-mail: rodriguez@minet.uni-jena.de

Runovskii Konstantin Mathematisches Institut, Fakultät für Mathematik und Informatik, Friedrich-Schiller-Universität, Leutragraben 1, 07743 Jena, Germany

Runst Thomas Mathematisches Institut, Fakultät für Mathematik und Informatik, Friedrich-Schiller-Universität, Leutragraben 1, 07743 Jena, Germany

E-mail: runst@minet.uni-jena.de

Sheremeta Myroslav Mykolayovych 700-richcha st. 99/55, 290020 Lviv, Ukraine

E-mail: kond@franko.lviv.ua

Schmeisser Hans-Jürgen Mathematisches Institut, Fakultät für Mathematik und Informatik, Friedrich-Schiller-Universität, Leutragraben 1, 07743 Jena, Germany

E-mail: schmeiszer@minet.uni-jena.de

Schott Thomas Mathematisches Institut, Fakultät für Mathematik und Informatik, Friedrich-Schiller-Universität, Leutragraben 1, 07743 Jena, Germany

E-mail: schott@minet.uni-jena.de

Sickel Winfried Mathematisches Institut, Fakultät für Mathematik und Informatik, Friedrich-Schiller-Universität, Leutragraben 1, 07743 Jena, Germany

E-mail: sickel@minet.uni-jena.de

Tarasyuk Svyatoslav Ivanovych Shevchenka st. 26/29, Dublyany, 292040 Zovkivskii, Lviv reg., Ukraine

E-mail: kond@franko.lviv.ua

Verdiev Timur Verdievich Engels st. 52-10, Makhachkhala, Russia Verdiev Verdi Gadzhimagomedovich Engels st. 52-10, Makhachkhala,

Russia Zabolotskii Mykola Vasylyovych Holovatskoho st. 17a/6,290018 Lviv,

Ukraine

E-mail: kond@franko.lviv.ua